solve log(2x+3) = log(4x)+ 2, for x giving your answer correct to 3 significant figures
Answer by fractalier(2101)

Provided the problem is correctly stated we have
log(2x+3) = log(4x) + 2
log(2x+3) - log(4x) = 2
log [(2x+3) / 4x] = 2
so that when we exponentiate we get
(2x+3) / 4x = 10^2 = 100
2x + 3 = 400x
3 = 398x
x = 3/398 or about .00754